# Approximate Nearest Neighbor Search with Window Filters

### v1

@@ 5.2 Analysis of Super Postfiltering
Intuitively, the blowup for $[a,b]$ is the ratio between the size of $[a,b]$ and the smallest range in $R$ that contains it. Note that if no range $[a_{i},b_{i}]$ contains $[a,b]$, we say that the blowup is equal to $\infty$. We can further define the worst case blowup for a set of ranges (like $R$) by taking the maximum blowup over all possible query ranges $[a,b]$:
@@
Intuitively, if we build an ANN index for the points corresponding to each range, then the worst-case blowup limits how expensive a query can be, while the $\mathsf{cost}$ approximates the memory required (since most practical ANN indices, e.g., LSH (Andoni & Indyk, 2008) and DiskANN (Jayaram Subramanya et al., 2019), have memory that is approximately linear in the number of points they index). Note that here, a range of, e.g., $[17,35]$ corresponds to an ANN index built on the $17$’th point through the $35$’th point in $D$, assuming the points are sorted by label value.
@@
We can construct an $R$ consisting of the ranges corresponding to each subset indexed in a $\beta$-WST, so we can analyze it using Definition 5.4.
@@
Table 2: Summary of datasets used in our experiments.
6 Experiments
Experiment Setup. We run all query methods on all datasets and filter widths on a 2.20GHz Intel Xeon machine with $40$ cores and two-way hyper-threading, $100$ MiB L3 cache, and $504$ GB of RAM. We run index building using all $80$ hyper-threads and restrict queries to $16$ threads, and parallelize across (and not within) queries. We run index construction experiments with varying $\beta$ on a separate 2.10GHz 4 socket Intel Xeon machine with $96$ cores and two way hyperthreading, $132$ MiB L3 cache, and $1.47$ TB of RAM. We use all threads on the machine for these experiments.
[…]
We note that our theory focuses on the $c$-ANN problem, which only concerns whether a single $c$-approximate nearest neighbor is returned. However, as is standard in much of the ANN literature, in our experiments we report recall of the top $10$ filtered neighbors to the query. While our runtime proofs in Theorem 5.1, Lemma 5.2, and Lemma 5.3 assume that the underlying ANNS algorithm returns a single ANN, in practice, we find that DiskANN has high recall for both single ANN and top-$10$ ANN. Therefore, we believe that our theoretical analyses still provide insights into our empirical results for top-$10$ ANN.
Finally, we ensure that our smallest filter fractions are still wide enough such that there are at least $10$ points that meet the filter constraint, i.e., we ensure $|D_{(a,b)}|\geq 10$.
Implementation Details. We provide a performant, open source C++ library and associated Python bindings.  [^1] https://github.com/JoshEngels/RangeFilteredANN Our code is built on the ParlayLib library (Blelloch et al., 2020) for parallel programming and the recent ParlayANN suite of parallel ANN algorithms (Dobson et al., 2023). We implement a number of memory and performance optimizations, including using a larger base case of $1000$ in Algorithm 1 and storing the entire dataset just once across all sub-indices.
Datasets. The datasets we use are listed in Table 2 and further explained below. All datasets are available in the same repository as the code and have free for research licenses; more licensing information is included in Section D.1.
@@ 6 Experiments
  • SIFT, GloVe, and Deep are ANN datasets from the widely used and standardized ANN benchmarks repository (Aumüller et al., 2020). To adapt them to window search, we uniformly at random generate a label for each point between 00 and $1$. We create $16$ different query sets $Q_{1},\ldots,Q_{16}$, each one using the same $10000$ query vectors from ANN benchmarks with random filters of fraction $2^{-i}$.
  • Adverse is a synthetic dataset tailored to disadvantage methods that rely on the label and point distributions being independent. The overall idea is to craft a dataset and queries where the points that meet the filter constraint are much farther away from the query than the rest of the dataset. To do this, we let $D$ be a mixture of $100$ Gaussians and draw $10000$ points from each Gaussian, where the means $\mu_{i}$ are drawn from $N(0,I)$ and each Gaussian is distributed as $N(\mu_{i},0.01\cdot I)$ ($I$ is the $100$ dimensional identity matrix). A point generated from the $i$’th Gaussian has a label equal to $i+\mathsf{Uniform}(-0.5,0.5)$, and we generate a query for every pair $i,j\in\{1,\ldots,100\}$ with $i\neq j$ that consists of a random point drawn from Gaussian $i$ with window filter $(j-0.5,j+0.5)$. In other words, each query filters to only points from a different cluster than it itself is drawn from.
Query Methods and Hyperparameters. We run all of the query methods described in Section 4.1, Section 4.2, and Section 4.3. For all methods that use an arbitrary index $A$, we use a DiskANN index as described earlier. We run Algorithm 2 with $\beta=2$ unless specified otherwise; we call this method $\mathsf{DiskANNWST}$ in our experiments. We use $\mathsf{Prefiltering}$ unmodified as described in Section 4.1. We expect $\mathsf{Prefiltering}$ to always achieve near 100% recall; it may not be 100% exactly due to numerical precision issues. For $\mathsf{Postfiltering}$, $\mathsf{OptimizedPostfiltering}$, $\mathsf{ThreeSplit}$, and $\mathsf{SuperPostfiltering}$ we search over initial $k$ in $[10,20,40,80,160,320,640,1280]$ and $\mathsf{final\_multiply}$ in $[1,2,3,4,8,16,32]$. We use the setting $\gamma=2$ from Theorem 5.6 for $\mathsf{SuperPostfiltering}$, which guarantees a worst case blowup factor of $4$ using in practice about $1.5$ times as much memory as $\mathsf{DiskANNWST}$.
$\mathsf{Milvus}$ treats window search as an instantiation of categorial filter search. Before querying the underlying ANN index, $\mathsf{Milvus}$ creates a bitset that marks all of the points that meet the window filter. Then, while traversing the underlying ANN index, Milvus ignores points that are not set in the bitset. We try all supported underlying $\mathsf{Milvus}$ indices: HNSW, IVF_PQ, IVF_SQ8, SCANN, and IVF_FLAT. We search over the same beam sizes as for $\mathsf{Postfiltering}$. $\mathsf{Milvus}$ does not natively support a batch query with different filters for each point in the batch, so we wrote a Python multiprocessing program that spins up many processes that query the constructed index in parallel.
In addition to the tricks described in Section 2, $\mathsf{VBASE}$ uses a query planning step to attempt to predict how many initial results to retrieve, then filters these retrieved results, before finally applying a final top-$k$ ranking to the retrieved results that meet the filter constraint. We were not able to get multiple queries to run in parallel with VBASE (and in the original paper they also only operate in the regime of a single query at a time).
@@
Figure 3: Comparison of window search methods on Adverse. Up and to the right is better. $\mathsf{DiskANNWST}$ and $\mathsf{ThreeSplit}$ achieve a good recall vs. latency tradeoff, but besides the $\mathsf{Prefiltering}$ baseline all of the other methods are unable to achieve a reasonable recall. All methods are run with $16$ threads.
@@
Table 3: Speedup of our best method over the best baseline, restricted to hyperparameter settings that yield at least $0.95$ recall. All methods are run with $16$ threads. We show a speedup across all filter fractions smaller than $2^{-1}$ on all datasets. We show up to a 75×75$\times$75 × speedup on medium filter fractions.
Tradeoff of Queries per Second vs. Recall. We plot the Pareto frontier of recall vs. queries per second of all methods on a selection of filter fractions on deep in Figure 2. We include plots of other datasets in the appendix; the main observations are substantially the same across datasets. Overall, our methods achieve multiple orders of magnitude query speedup at fixed recall levels. $\mathsf{SuperPostfiltering}$ does particularly well at lower recall levels of around $0.9$ to $0.99$, while at high recall levels $\mathsf{DiskANNWST}$ and $\mathsf{ThreeSplit}$ are the best methods, attaining about the same performance. Additionally, for small filter fractions the $\mathsf{Prefiltering}$ baseline is competitive with our methods, whereas for large filter fractions the $\mathsf{Postfiltering}$ baseline is competitive. Our methods achieve the largest speedups over baselines in the medium filter fractions. We note that among our methods, $\mathsf{OptimizedPostfiltering}$ performs the worst, which we explain with the high worst case blowup of a $2$-WST (see Lemma 5.5). Finally, we note that the two vector databases that we test against, $\mathsf{VBASE}$ and $\mathsf{Milvus}$, are completely dominated by our naive baselines $\mathsf{Prefiltering}$ and $\mathsf{Postfiltering}$.
We also plot results on Adverse in Figure 3. $\mathsf{Prefiltering}$ does well, as expected. Furthermore, $\mathsf{DiskANNWST}$ offers a good tradeoff between recall and query latency, which makes sense because the query time guarantee from Lemma 5.3 (assuming DiskANN also does well for top-$10$ ANN) holds for any query distribution, even an adversarial one. However, all of the methods that rely solely on postfiltering, as well as the other baselines, fail to achieve meaningful recall. For the postfiltering methods, this result is unsurprising: the index that gets selected for postfiltering has some points that do not meet the query’s filter constraints, and by construction these points are frequently closer to the query than the entire target cluster. Surprisingly, although $\mathsf{ThreeSplit}$ does use postfiltering as a subroutine, it is able to achieve similar performance to $\mathsf{DiskANNWST}$; this may be because after querying for the indices that make up the center of the label range (which is not postfiltering), the label ranges that are left on each side are typically much smaller.
We also include speedups of our best method (the best of $\mathsf{DiskANNWST}$, $\mathsf{OptimizedPostfiltering}$, $\mathsf{SuperPostfiltering}$, and $\mathsf{ThreeSplit}$) over the best baseline method on filter fractions $i=2^{0},\ldots,2^{-11}$ on all datasets in Table 3 at a recall level of $0.95$, and we include the same table at additional recalls in the appendix. These reinforce our findings in Figure 2 across other datasets; at a $0.95$ recall level, we achieve up to a $75$X speedup on Deep, up to a $16$X speedup on SIFT, up to a $9$X speedup on GloVe, and up to a $17$X speedup on Redcaps.
Dataset DiskANN $2$-WST Super
SIFT 1 m 7.5 m 13.5 m
GloVe 2.5 m 16.5 m 28 m
Deep 16.5 m 2 h 4 h
Redcaps 1.5 h 7 h 18.5 h
Adverse 2.5 m 23 m 41 m
Table 4: Build times for different indexing methods across all datasets, rounded to the nearest half unit. Index construction was done using $80$ threads.
Dataset Raw DiskANN $2$-WST Super
@@
Index Memory and Construction Time. Table 4 and Table 5 show the construction time and memory sizes for a single DiskANN index (for $\mathsf{Postfiltering}$), a $2$-WST with DiskANN indices at each node (for $\mathsf{DiskANNWST}$, $\mathsf{OptimizedPostfiltering}$, and $\mathsf{ThreeSplit}$), and the index for $\mathsf{SuperPostfiltering}$. We also show the memory for just the dataset, which is exactly how much memory $\mathsf{Prefiltering}$ needs (the build for $\mathsf{Prefiltering}$ is just a sort and takes less than a few seconds across all datasets). We see that the $2$-WST takes about 33338888 times as much memory as a single DiskANN index, depending on the dataset, whereas the index for $\mathsf{SuperPostfiltering}$ takes about 555514141414 times as much memory as a single DiskANN index. The construction times show a similar increase across methods, with a 555510101010X increase in build time from DiskANN to $2$-WST and a 1010101020202020X increase from DiskANN to $\mathsf{SuperPostfiltering}$. Finally, we note that while the larger datasets do take signficantly longer to build, we are using a large beam search construction buffer of $500$ in order to vary as few hyperparameters as possible, and a smaller buffer size may give faster build times with minimal loss in recall.
Figure 4: Pareto curves of recall vs. throughput on the SIFT dataset for a filter fraction of $2^{-1}$ and varied branching factors. Up and to the right is better. All trials are run with $16$ threads.
Varying $\beta$. A larger branching factor $\beta$ decreases the build time and memory footprint of a $\mathsf{DiskANNWST}$ index by reducing the number of levels in the tree and thus reducing the number of graphs that are built (see Figure 9 in the appendix for a plot showing the exact reduction in memory and build time as we scale $\beta$). However, as shown in Figure 4, this comes at the cost of a reduction in recall and latency. These results are substantially the same across different filter fractions; see Figure 10 in the appendix for experiments with more filter fractions.
@@ Appendix A Proofs
There are two components to proving Theorem 5.1. First, we must show that Algorithm 2 solves the $c$-approximate window search problem, and then we must show that it solves it in the given running time.
@@
$O\left(\log_{\beta}(N)*(\beta+\beta d)+(2\beta-2)*\sum_{j=0}^{\log_{\beta}(N)}A_{q}(D,N*\beta^{-j})\right)=O\left(\beta\log_{\beta}(N)d+\beta*\sum_{j=0}^{\log_{\beta}(N)}A_{q}(D,N*\beta^{-j})\right)$
See 5.2
@@
$=O\left(\frac{C\beta dN^{\rho}}{1-\beta^{-\rho}}\right)$
@@
$=O\left(\beta\log_{\beta}(N)(d+A_{q}(D))\right)$
See 5.3
@@
(Indyk & Xu, 2023) considers greedy search on an $\alpha$-DiskANN graph built on a dataset $D$ with ”slow preprocessing.” They show that the search procedure is guaranteed to return a $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN in $O(\log_{\alpha}(\frac{\Delta}{(\alpha-1)\epsilon}))$ steps, each of which take $O((4\alpha)^{\delta}\log\Delta)$. See Appendix C for an explanation of $\alpha$ and a complete overview of DiskANN.
@@
$O\left(\log_{\alpha}\left(\frac{\Delta}{\left(\alpha-1\right)\epsilon}\right)\left(4\alpha\right)^{\delta}\log\Delta\right)=O\left(\log_{\alpha}\left(\frac{\Delta}{\left(\alpha-1\right)(c-\frac{\alpha+1}{\alpha-1})}\right)\left(4\alpha\right)^{\delta}\log\Delta\right)$
@@
$\bigcup_{(j+2)m\leq N}\left\{[jm+1,(j+1)m+1]\right\}\bigcup[N-2m+1,N-m+1]\supset[1,N-2m]\bigcup[N-2m+1,N-m+1]=[1,N-m+1]$
@@
$=N\left(2\log_{\gamma}(N)+1\right)$

### v2

@@ 5.2 Analysis of Super Postfiltering
Intuitively, the blowup for $[a,b]$ is the ratio between the size of $[a,b]$ and the smallest range in $R$ that contains it. We can further define the worst case blowup for a set of ranges (like $R$) by taking the maximum blowup over all possible query ranges $[a,b]$:
@@
Intuitively, if we build an ANNS index for the points corresponding to each range, then the worst-case blowup limits how expensive a query can be, while the $\mathsf{cost}$ approximates the memory required (since most practical ANNS indices, e.g., LSH (Andoni & Indyk, 2008) and Vamana (Jayaram Subramanya et al., 2019), have memory that is approximately linear in the number of points that they index). Note that here, a range of, e.g., $[17,35]$ corresponds to an ANNS index built on the $17$’th point through the $35$’th point in $D$, assuming the points are sorted by label value.
@@
We can construct an $R$ consisting of the ranges corresponding to each subset indexed in a $\beta$-WST, so we can analyze it using Definition 5.5. We prove the following lemma.
@@
6 Experiments
Experiment Setup. We run all query methods on all datasets and filter widths on a 2.20GHz Intel Xeon machine with $40$ cores and two-way hyper-threading, $100$ MiB L3 cache, and $504$ GB of RAM. We run index building using all $80$ hyper-threads and restrict queries to $16$ threads, and parallelize across (and not within) queries. We run index construction experiments with varying values of $\beta$ on a separate 2.10GHz Intel Xeon machine with $96$ cores and two way hyper-threading, $132$ MiB L3 cache, and $1.47$ TB of RAM. We use all hyper-threads on the machine for these experiments.
We use a Vamana index (Jayaram Subramanya et al., 2019) with $\alpha=1$, $\textit{degree}{}=64$, and the construction beam search width set to $500$ for all ANNS indices, except for the $\mathsf{Milvus}$ and $\mathsf{VBASE}$ baselines. Vamana allows searching for $k\geq 1$ nearest neighbors; for simplicity of presentation, we assume that the query beam search width is always set to $k$. We defer a description of Vamana and its associated hyper-parameters to Appendix C.
We note that our theory focuses on the $c$-ANN problem, which only concerns whether a single $c$-approximate nearest neighbor is returned. However, as is standard in much of the ANN literature, in our experiments we report the recall of the top $10$ filtered neighbors to the query. While our runtime proofs in Theorem 5.2, Lemma 5.3, and Lemma 5.4 assume that the underlying ANNS algorithm returns a single ANN, in practice, we find that Vamana has high recall for both single ANN and top-$10$ ANN. Therefore, we believe that our theoretical analyses still provide insights into our empirical results for top-$10$ ANN.
Finally, we ensure that our smallest filter fractions are still wide enough such that there are at least $10$ points that meet the filter constraint, i.e., we ensure that $|D_{(a,b)}|\geq 10$.
Implementation Details. We provide an open source C++ library and associated Python bindings. [^1] https://github.com/JoshEngels/RangeFilteredANN Our code is built on the ParlayLib library (Blelloch et al., 2020) for parallel programming and the recent ParlayANN suite of parallel ANNS algorithms (Manohar et al., 2024). We implement a number of memory and performance optimizations, including using a larger base case of $1000$ in Algorithm 1 and storing the entire dataset just once across all sub-indices.
Table 2: Summary of datasets used in our experiments.
@@ 6 Experiments
Datasets. The datasets that we use are listed in Table 2 and further explained below. All datasets are available in the same repository as the code; licensing information is included in Section D.1.
  • SIFT, GloVe, and Deep are ANN datasets from the widely used and standardized ANN benchmarks repository (Aumüller et al., 2020). To adapt them to window search, we generate a label for each point uniformly at random between 00 and $1$. We create $16$ different query sets $Q_{1},\ldots,Q_{16}$, each one using the same $10000$ query vectors from ANN benchmarks with random filters of fraction $2^{-i}$.
  • Adverse is a synthetic dataset tailored to disadvantage methods that rely on the label and point distributions being independent. The overall idea is to craft a dataset and queries where the points that meet the filter constraint are much farther away from the query than the rest of the dataset. To do this, we let $D$ be a mixture of $100$ Gaussians and draw $10000$ points from each Gaussian, where the means $\mu_{i}$ are drawn from $N(0,I)$ and each Gaussian is distributed as $N(\mu_{i},0.01\cdot I)$ ($I$ is the $100$-dimensional identity matrix). A point generated from the $i$’th Gaussian has a label equal to $i+\mathsf{Uniform}(-0.5,0.5)$, and we generate a query for every pair $i,j\in\{1,\ldots,100\}$ with $i\neq j$ that consists of a random point drawn from Gaussian $i$ with window filter $(j-0.5,j+0.5)$. In other words, each query filters to only points from a different cluster than it itself is drawn from.
Query Methods and Hyper-parameters. We run all of the query methods described in Section 4.1, Section 4.2, and Section 4.3. For all methods that use an arbitrary index $A$, we use a Vamana index as described earlier. We run Algorithm 2 with $\beta=2$ unless specified otherwise; we call this method $\mathsf{VamanaWST}$ in our experiments. We use $\mathsf{Prefiltering}$ unmodified as described in Section 4.1. We expect $\mathsf{Prefiltering}$ to always achieve near 100% recall; it may not be 100% exactly due to numerical precision issues. For $\mathsf{Postfiltering}$, $\mathsf{OptimizedPostfiltering}$, $\mathsf{ThreeSplit}$, and $\mathsf{SuperPostfiltering}$, we search over initial values of $k$ in $[10,20,40,80,160,320,640,1280]$ and $\mathsf{final\_multiply}$ value in $[1,2,3,4,8,16,32]$. We use the setting $\gamma=2$ from Theorem 5.7 for $\mathsf{SuperPostfiltering}$, which guarantees a worst case blowup factor of $4$ using in practice about $1.5$ times as much memory as $\mathsf{VamanaWST}$.
$\mathsf{Milvus}$ treats window search as an instantiation of categorial filter search. Before querying the underlying ANNS index, $\mathsf{Milvus}$ creates a bitset that marks all of the points that meet the window filter. Then, while traversing the underlying ANNS index, Milvus ignores points that are not set in the bitset. We try all supported underlying $\mathsf{Milvus}$ indices: HNSW, IVF_PQ, IVF_SQ8, SCANN, and IVF_FLAT. We search over the same beam sizes as for $\mathsf{Postfiltering}$. $\mathsf{Milvus}$ does not natively support a batch query with different filters for each point in the batch, so we wrote a Python multiprocessing program that spins up many processes that query the constructed index in parallel.
In addition to the heuristic described in Section 2, $\mathsf{VBASE}$ uses a query planning step to attempt to predict how many initial results to retrieve, then filters these retrieved results, before finally applying a final top-$k$ ranking to the retrieved results that meet the filter constraint. We were not able to get multiple queries to run in parallel with VBASE (and in the original paper they also only operate in the regime of a single query at a time).
@@
Tradeoff of Queries per Second vs. Recall. We plot the Pareto frontier of recall vs. queries per second of all methods on a selection of filter fractions on Deep in Figure 2 (the indices we compute the frontier on correspond to the hyper-parameters specified in the last section). We include plots of other datasets in Appendix D; the main observations are substantially the same across datasets. Overall, our methods achieve multiple orders of magnitude query speedup at fixed recall levels. $\mathsf{SuperPostfiltering}$ does particularly well at lower recall levels of around $0.9$ to $0.99$, while at high recall levels $\mathsf{VamanaWST}$ and $\mathsf{ThreeSplit}$ are the best methods, attaining about the same performance. Additionally, for small filter fractions the $\mathsf{Prefiltering}$ baseline is competitive with our methods, whereas for large filter fractions the $\mathsf{Postfiltering}$ baseline is competitive. Our methods achieve the largest speedups over baselines in the medium filter fractions. We note that among our methods, $\mathsf{OptimizedPostfiltering}$ performs the worst, which we explain with the high worst case blowup of a $2$-WST (see Lemma 5.6). Finally, we note that the two vector databases that we test against, $\mathsf{VBASE}$ and $\mathsf{Milvus}$, are completely dominated by our naive baselines $\mathsf{Prefiltering}$ and $\mathsf{Postfiltering}$.
@@
Figure 3: Comparison of window search methods on Adverse. Up and to the right is better. $\mathsf{VamanaWST}$ and $\mathsf{ThreeSplit}$ achieve a good recall vs. latency tradeoff, but besides the $\mathsf{Prefiltering}$ baseline all of the other methods are unable to achieve a reasonable recall. All methods are run with $16$ threads.
We also plot results on Adverse in Figure 3. $\mathsf{Prefiltering}$ does well, as expected. Furthermore, $\mathsf{VamanaWST}$ offers a good tradeoff between recall and query latency, which makes sense because the query time guarantee from Lemma 5.4 (assuming Vamana also does well for top-$10$ ANN) holds for any query distribution, even an adversarial one. However, all of the methods that rely solely on postfiltering, as well as the other baselines, fail to achieve meaningful recall. For the postfiltering methods, this result is unsurprising: the index that gets selected for postfiltering has many points that do not meet the query’s filter constraints, and by construction these points are frequently closer to the query than the entire target cluster. The beam search then has to expand over many near neighbors to reach points meeting the filter constraint, at which point the accuracy of the beam search is likely low. Surprisingly, although $\mathsf{ThreeSplit}$ does use postfiltering as a subroutine, it is able to achieve similar performance to $\mathsf{VamanaWST}$; this may be because after querying for the indices that make up the center of the label range (which is not postfiltering), the label ranges that are left on each side are typically much smaller.
Table 3: Speedup of our best method over the best baseline, restricted to hyper-parameter settings that yield at least $0.95$ recall. All methods are run with $16$ threads. We show a speedup across all filter fractions smaller than $2^{-1}$ on all datasets. We show up to a 75×75$\times$75 × speedup on medium filter fractions.
@@
We also include speedups of our best method (the best of $\mathsf{VamanaWST}$, $\mathsf{OptimizedPostfiltering}$, $\mathsf{SuperPostfiltering}$, and $\mathsf{ThreeSplit}$) over the best baseline method on filter fractions $i=2^{0},\ldots,2^{-11}$ on all datasets in Table 3 at a recall level of $0.95$, and we include the same table at additional recall levels in Appendix D. These reinforce our findings in Figure 2 across other datasets; at a $0.95$ recall level, we achieve up to a $75$X speedup on Deep, up to a $16$X speedup on SIFT, up to a $9$X speedup on GloVe, and up to a $17$X speedup on Redcaps.
Table 4: Build times for different indexing methods across all datasets, rounded to the nearest unit. Index construction was done using $80$ threads.
Dataset Vamana $2$-WST Super
SIFT 1 m 8 m 14 m
GloVe 3 m 17 m 28 m
Deep 17 m 2 h 4 h
Redcaps 2 h 7 h 19 h
Adverse 3 m 23 m 41 m
Dataset Raw Vamana $2$-WST Super
@@
Index Memory and Construction Time. Table 4 and Table 5 show the construction time and memory sizes for a single Vamana index (for $\mathsf{Postfiltering}$), a $2$-WST with Vamana indices at each node (for $\mathsf{VamanaWST}$, $\mathsf{OptimizedPostfiltering}$, and $\mathsf{ThreeSplit}$), and the index for $\mathsf{SuperPostfiltering}$. We also show the memory for just the dataset, which is exactly how much memory $\mathsf{Prefiltering}$ needs (the build for $\mathsf{Prefiltering}$ is just a sort and takes less than a few seconds across all datasets). We see that the $2$-WST takes about 33338888 times as much memory as a single Vamana index, depending on the dataset, whereas the index for $\mathsf{SuperPostfiltering}$ takes about 555514141414 times as much memory as a single Vamana index. The construction times show a similar increase across methods, with a 555510101010X increase in build time from Vamana to $2$-WST and a 1010101020202020X increase from Vamana to $\mathsf{SuperPostfiltering}$. Finally, we note that while the larger datasets do take signficantly longer to build, we are using a large beam search construction buffer of $500$ in order to vary as few hyper-parameters as possible, and a smaller buffer size may give faster build times with minimal loss in recall.
Figure 4: Pareto curves of recall vs. throughput on the SIFT dataset for a filter fraction of $2^{-1}$ and varying branching factors $\beta$ for $\mathsf{VamanaWST}$. Up and to the right is better. All trials are run with $16$ threads.
Varying $\beta$. A larger branching factor $\beta$ decreases the build time and memory footprint of a $\mathsf{VamanaWST}$ index by reducing the number of levels in the tree and thus reducing the number of graphs that are built (see Figure 9 in Appendix D for a plot showing the exact reduction in memory and build time as we scale $\beta$). However, as shown in Figure 4, this comes at the cost of a reduction in recall and latency. These results are substantially the same across different filter fractions; see Figure 10 in Appendix D for experiments with more filter fractions.
@@ Appendix A Proofs
As stated in the main text, if we have some function parameterized by the dataset and subset size $O(A_{f}(D,m))$, then this function evaluated on all nodes of Algorithm 1 is
$O\left(\sum_{j=0}^{\log_{\beta}N}\beta^{j}A_{f}(D,N\cdot\beta^{-j})\right).$ (1)
If $A_{f}$ is of the form $Cm^{\rho}$ for $\rho\geq 1$ and for some constant $C$ depending on $D$, then this is equivalent to:
$=O\left(CN^{\rho}\sum_{j=0}^{\log_{\beta}N}\beta^{j-j\rho}\right)$
$=O\left(CN^{\rho}\sum_{j=0}^{\log_{\beta}N}(\beta^{1-\rho})^{j}\right)$
=$=\begin{cases}O(CN^{\rho}\log_{\beta}N)&\text{if $\rho=1$}\\ O\left(CN^{\rho}\frac{1-N^{1-\rho}}{1-\beta^{1-\rho}}\right)=O((\frac{1}{1-\beta^{1-\rho}})CN^{\rho})&\text{if $\rho>1$}\end{cases}$
Determining the memory of the index returned by Algorithm 1 is equivalent to determining the memory of all nodes built on Line 5 throughout the recursion. Similarly, as long as these calls take longer than constant time, they are the computational bottleneck of the recursion. Thus, we can plug the Vamana build time and memory from (Indyk & Xu, 2023) into these results to get the build time and memory of a Vamana WST.
The slow preprocessing version of Vamana takes $O(N^{3})$ for construction time and takes up $O(N(4\alpha)^{\delta}\log\Delta)$ space, so plugging into these results we have that a $\beta$-WST tree with a slow preprocessing Vamana implementation takes $O(\frac{1}{1-\beta^{-2}}N^{3})=O(N^{3})$ time to build and has memory size $O((4\alpha)^{\delta}\log(\Delta)N\log_{\beta}N)$. ∎
See 5.2
Proof.
First, we will show that Algorithm 2 solves the $c$-approximate window search problem. Then, we will show that it solves it in the given running time.
@@
$O\left(\log_{\beta}(N)*(\beta+\beta d)+(2\beta-2)*\sum_{j=0}^{\log_{\beta}(N)}A_{q}(D,N*\beta^{-j})\right)=O\left(\beta\log_{\beta}(N)d+\beta*\sum_{j=0}^{\log_{\beta}(N)}A_{q}(D,N*\beta^{-j})\right).$
See 5.3
@@
$=O\left(\frac{C\beta dN^{\rho}}{1-\beta^{-\rho}}\right).$
@@
$=O\left(\beta\log_{\beta}(N)(d+A_{q}(D))\right).$
See 5.4
@@
(Indyk & Xu, 2023) consider greedy search on an $\alpha$-Vamana graph built on a dataset $D$ with ”slow preprocessing.” They show that the search procedure is guaranteed to return a $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN in $O(\log_{\alpha}(\frac{\Delta}{(\alpha-1)\epsilon}))$ steps, each of which take $O((4\alpha)^{\delta}\log\Delta)$ time. See Appendix C for an explanation of $\alpha$ and a complete overview of Vamana.
@@
$O\left(\log_{\alpha}\left(\frac{\Delta}{\left(\alpha-1\right)\epsilon}\right)\left(4\alpha\right)^{\delta}\log\Delta\right)=O\left(\log_{\alpha}\left(\frac{\Delta}{\left(\alpha-1\right)(c-\frac{\alpha+1}{\alpha-1})}\right)\left(4\alpha\right)^{\delta}\log\Delta\right).$
@@
$\bigcup_{(j+2)m\leq N}\left\{[jm+1,(j+1)m+1]\right\}\bigcup[N-2m+1,N-m+1]\supset[1,N-2m]\bigcup[N-2m+1,N-m+1]=[1,N-m+1].$
@@
$=N\left(2\log_{\gamma}(N)+1\right).$